OpenStudy (calculusxy):

MEDAL FOR YOUR HELP Maria uses 5 1/3 gallons of gas to drive to work and back four times. a. How many gallons of gas does Maria use in one round trip to work? b. Maria's car gets 28 miles to the gallon. How many miles is her round trip to work?

OpenStudy (anonymous):

A. Number of gallons per round trip = gas used / number of round trips = 5 1/3 ÷ 4 = 16/3 x 1/4 [dividing by 4 is the same as multiplying by 1/4] = 4/3 = 1 1/3 gallons per round trip

OpenStudy (calculusxy):

You got that from Yahoo Answers. Right?

OpenStudy (anonymous):

B. Number of miles per round trip = 1 1/3 gallons x 28 miles per gallon = 4/3 x 28/1 = (4 x 28) / 3 = 112/3 = 37 1/3 miles per round trip

OpenStudy (anonymous):

yea i figured it be easier than post in the link

OpenStudy (calculusxy):

Thanks :)

OpenStudy (anonymous):

mhm c;